# Hölder curves with exotic tangent spaces

### v2

@@ 8.1. Local cut-points in a class of carpets
To verify (i), fix such an open set $V\subset\mathbb{R}^{2}$. Then there exist $z\in V$ and $\epsilon>0$ such that $B(z,\epsilon)\subset V$ and $B(z,\epsilon/2)\cap L_{k}\neq\emptyset$. Assume for the sake of contradiction that (i) fails to hold for this $V$, and thus $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$ infinitely often. Now by Lemma 8.10 and since $\lim_{j\to\infty}Y_{N_{j}}=L_{k}$, there exists $j_{0}\in\mathbb{N}$ so that for every integer $j\geq j_{0}$, $\operatorname{excess}(Y_{N_{j}}\cap\overline{B}({\bf 0},|z|+\epsilon),X_{N_{j}})<\epsilon/2$ and $B(z,\epsilon/2)\cap Y_{N_{j}}\neq\emptyset$. By our assumption, there exists an integer $j\geq j_{0}$ such that $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$ and there exists a point $y\in Y_{N_{j}}\cap B(z,\epsilon/2)$. Then we have
@@
Now to verify (ii), fix $\delta>0$ and $z\in\mathbb{R}^{2}$ so that $B(z,\delta)\cap L_{k}=\emptyset$ and let $\epsilon\in(0,\delta)$. Since $Y_{N_{j}}\to L_{k}$ as $j\to\infty$ with respect to the Attouch-Wets topology, there is a $j_{0}\in\mathbb{N}$ so that for every integer $j\geq j_{0}$, $Y_{N_{j}}\cap B(z,\frac{\epsilon+\delta}{2})=\emptyset$. Furthermore by Lemma 8.10, choosing $j_{0}$ large enough, we may assume that for every integer $j\geq j_{0}$, $\operatorname{excess}(X_{N_{j}}\cap\overline{B}({\bf 0},|z|+\delta),Y_{N_{j}})<\frac{\delta-\epsilon}{2}$. Then for every integer $j\geq j_{0}$ and for each $p\in X_{N_{j}}\cap B({\bf 0},|z|+\delta)$, $\operatorname{dist}(p,Y_{N_{j}})<\frac{\delta-\epsilon}{2}$. Since $Y_{N_{j}}\cap B(z,\frac{\epsilon+\delta}{2})=\emptyset$, if $p\in B(z,\epsilon)$, then $\operatorname{dist}(p,Y_{N_{j}})\geq\frac{\delta-\epsilon}{2}$ and $p\in B({\bf 0},|z|+\delta)$. Therefore for each integer $j\geq j_{0}$, $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$, as desired. ∎
@@
Assume first that $t\geq 1$. Let $x_{0}\in X$, let $T\in\operatorname{Tan}(X,x_{0})$, and let $r_{j}\to 0$ be a sequence of scales so that $r_{j}^{-1}(X-x_{0})\to T$ as $j\to\infty$ in the Attouch-Wets topology. For each $j\in\mathbb{N}$, define the function
@@
Let $R>0$, let $\epsilon>0$, and let $j_{0}\in\mathbb{N}$ so that for every integer $j\geq j_{0}$,
@@
The other excess estimate follows via a similar calculation, so $r_{j}^{-1/t}(Y-f(x_{0}))\to F(T)$ as $j\to\infty$ in the Attouch-Wets topology. This concludes the case $t\geq 1$.
Now let $t\in(0,1)$. Then $f^{-1}:Y\to X$ is a $t$-bi-Hölder homeomorphism. Fix $x_{0}\in X$, let $T\in\operatorname{Tan}(X,x_{0})$, and let $r_{j}\to 0$ be a sequence of scales so that $r_{j}^{-1}(X-x_{0})\to T$ as $j\to\infty$ in the Attouch-Wets topology. Then $r_{j}^{t}\to 0$ is a sequence of scales, and there is a convergent subsequence $r_{j_{k}}^{-t}(Y-f(x_{0}))\to S$ as $k\to\infty$ in the Attouch-Wets topology. As in the case $t\geq 1$, we have $g_{j_{k}}:r_{j_{k}}^{-t}(Y-f(x_{0}))\to r_{j_{k}}^{-1}(X-x_{0})$, a sequence of $t$-bi-Hölder homeomorphisms with uniform coefficients which extend to $t$-Hölder functions $G_{j_{k}}:\mathbb{R}^{m}\to\mathbb{R}^{d}$, and the maps $G_{j_{k}}$ converge locally uniformly to a $t$-Hölder map $G:\mathbb{R}^{m}\to\mathbb{R}^{d}$. Then
@@
By Assouad’s embedding theorem [Ass83], there exists $N\in\mathbb{N}$ and a bi-Lipschitz embedding of the metric space $(K^{\eta},|\cdot|^{\alpha_{n}/s})$ into $\mathbb{R}^{N}$ and denote by $\Gamma$ the embedded image. This mapping produces a $(\alpha_{n}/s)$-bi-Hölder homeomorphism $g:K^{\eta}\to\Gamma$. Therefore, there exists some $C>1$ such that for every Borel $B\subset K^{\eta}$,
@@
so $\mathcal{H}^{s}(\Gamma)>0$. Moreover, the map $g\circ f:[0,1]\to\Gamma$ is a $(1/s)$-Hölder surjection. By Lemma 9.1, we have that for every point $x\in K^{\eta}$ and every tangent $T\in\operatorname{Tan}(K^{\eta},x)$ there exists a tangent $T^{\prime}\in\operatorname{Tan}(\Gamma,g(x))$ so that $T$ and $T^{\prime}$ are $(\alpha_{n}/s)$-bi-Hölder equivalent; in particular, $T$ and $T^{\prime}$ are homeomorphic to each other. Therefore, $\operatorname{Tan}(\Gamma,x)$ contains infinitely many topologically distinct elements. ∎
@@ 10. Tangents of self-similar sets
In this section we prove Proposition 1.3 and Theorem 1.4. We start with the proof of Proposition 1.3.
Proof of Proposition 1.3.
@@
Fix $x\in Q\cap U$, $S\in\operatorname{Tan}(Q,x)$, and $R>0$. Let $r_{j}\to 0$ be a decreasing sequence of positive scales so that $r_{1}<1/R$ and the sets $D_{j}:=(r_{j})^{-1}(Q-x)$ converge to $S$ in the Attouch-Wets topology as $j\to\infty$. Define for each $j\in\mathbb{N}$ the similarity $g_{j}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, given by $g_{j}(y)=(r_{j})^{-1}(y-x)$, and note that $g_{j}(Q)=D_{j}$. Define the alphabet $A=\{1,\dots,m\}$ and for a word $w\in A^{*}$, define $L_{w}:=\operatorname{Lip}(\phi_{w})$. Following [BV21, §2.3], for any $\delta\in(0,1)$ define
@@
Note that for each $j$, $\overline{B}(x,r_{j}R)\cap Q\subset\bigcup_{w\in C^{R}_{j}}\phi_{w}(Q)$. Additionally, by the OSC, if $v,w\in A^{*}$ are distinct words so that $\phi_{w}(U)\cap\phi_{v}(U)\neq\emptyset$, then there exists $u\in A^{*}\setminus\{\varepsilon\}$ so that either $w=vu$ or $v=wu$, which further implies that $\frac{L_{w}}{L_{v}}\leq L_{1}$ or $\frac{L_{v}}{L_{w}}\leq L_{1}$. Moreover, if $j\in\mathbb{N}$ and $w,v\in C^{R}_{j}$, then $\frac{L_{w}}{L_{v}}>L_{1}$ (or vice-versa); thus if $w,v\in C^{R}_{j}$ are distinct, $\phi_{w}(U)\cap\phi_{v}(U)=\emptyset$. Roughly speaking, each set $\{\phi_{w}:w\in C^{R}_{j}\}$ respects the “disjoint images” part of the open set condition for $U$. Observe that for each $j\in\mathbb{N}$ and each $w\in C^{R}_{j}$ by (10.1), the fact that $\operatorname{diam}(Q)=1$ and the fact that $\phi_{w}(Q)\cap\overline{B}(x,r_{j}R)\neq\emptyset$, we have $\phi_{w}(Q)\subset\overline{B}(x,2r_{j}R)\cap Q$. Hence, for each $j\in\mathbb{N}$,
@@
Next, since $Q$ is self-similar, it follows that $\mathcal{H}^{s}(Q)>0$ [Hut81, §5.3] and, in fact, $Q$ is Ahlfors $s$-regular [BV21, Lemma 2.4]. That is, there exists $c_{1}\geq 1$ so that for every $y\in Q$ and $r\in(0,1)$,
@@
Furthermore, if $\{f_{1}^{2R},\dots,f_{M_{2R}}^{2R}\}$ is constructed in a similar manner (replacing $R$ by $2R$), then for every $i\in\{1,\dots,M_{R}\}$ there exists unique $j\in\{1,\dots,M_{2R}\}$ so that $f_{i}^{R}(Q)\subset f_{j}^{2R}(Q)$. Indeed, the sets $C_{j}^{R}$ and $C_{j}^{2R}$ satisfy $\bigcup_{v\in C_{j}^{R}}\phi_{v}(Q)\subset\bigcup_{u\in C_{j}^{2R}}\phi_{u}(Q)$, so for each $v\in C_{j}^{R}$, there exists a unique $u\in C_{j}^{2R}$ so that $\phi_{v}(U\cap Q)\subset\phi_{u}(U\cap Q)$. Then it must hold that $f_{i}^{R}(U\cap Q)$ is an open subset of $f_{j}^{2R}(U\cap Q)\cap B({\bf 0},2R)$ which is an open subset of $S\cap B({\bf 0},2R)$. Thus, $f_{i}^{R}(U\cap Q)$ is an open subset of $S$. ∎
The following remark follows easily from the proof of Proposition 1.3.
Remark 10.1.
Let $Q$ and $S$ be as in Proposition 1.3. Then the Hausdorff dimensions of $Q$ and of $S$ are equal and they satisfy $\mathcal{H}^{\operatorname{dim_{H}}(Q)}(S)>0$, $\mathcal{H}^{\operatorname{dim_{H}}(Q)}(Q)\in(0,\infty)$.
10.1. Self-similar sponges
(10.3) $\{S_{i}(y)=k^{-1}y+p_{i}:[0,1]^{N}\to[0,1]^{N}\}_{i=1}^{m},$
Lemma 10.2.
Proof.
$e:=S_{1}({\bf 0})-S_{2}({\bf 0})\in\{-k^{-1},0,k^{-1}\}^{N}.$
$S_{1}([0,1]^{N})\cap S_{2}([0,1]^{N})=S_{1}(I_{1}\times\dots\times I_{N}).$
For the case $l\geq 3$, we have that
@@ 10.1. Self-similar sponges
Suppose that $K$ is the attractor of $\{S_{i}\}_{i=1}^{m}$ with $S_{i}(y)=k^{-1}y+p_{i}$ where $k$ and $\{p_{1},\dots,p_{m}\}$ are as in (10.3). If $m=1$, then the result is trivial, so we may assume that $m\geq 2$. Set $A=\{1,\dots,m\}$. Note that for each $n\in\mathbb{N}$, $v\in A^{n}$, and $i\in A$,
@@
If $(0,1)^{N}\cap K=\emptyset$, then the result is trivial, so assume $(0,1)^{N}\cap K\neq\emptyset$. Fix $x\in(0,1)^{N}\cap K$, $T\in\operatorname{Tan}(K,x)$, $p,q\in T$, $\rho_{1},\rho_{2}>0$, and a sequence of positive scales $r_{j}\to 0$ so that the sets $X_{j}:=(r_{j})^{-1}(K-x)$ converge to $T$ in the Attouch-Wets topology as $j\to\infty$.
@@
We claim that $\bigcap_{w\in W_{j}}S_{w}([0,1]^{N})\neq\emptyset$ for every $j\in\mathbb{N}$. We first show the claim for two words only. Fix distinct $v,w\in W_{j}$ and let $z_{v},z_{w}\in[0,1]^{N}$ so that
@@
so by the triangle inequality we obtain
@@
By Lemma 10.2, for every $j\in\mathbb{N}$ there exists a face $C_{j}$ of $0,1]^{N}$ so that $\bigcap_{i=1}^{M}S_{v_{i}^{j}}([0,1]^{N})=S_{v_{1}^{j}}(C_{j})$. Thus, for every $j\in\mathbb{N}$ there exist distinct $e_{j,1},\dots,e_{j,M}\in\{0,1\}^{N}$ so that $S_{v_{i}^{j}}(e_{j,i})=S_{v_{1}^{j}}(e_{j,1})$ for every $i\in\{1,\dots,M\}$. Passing to a subsequence, we may assume that there exist distinct $e_{1},\dots,e_{M}\in\{0,1\}^{N}$ so that $e_{j,i}=e_{i}$ for every $j\in\mathbb{N}$ and $i\in\{1,\dots,M\}$. Then $f_{i}(e_{i})=f_{1}(e_{1})$ and $S_{v_{i}^{j}}(e_{i})=S_{v_{1}^{j}}(e_{1})$ for every $j\in\mathbb{N}$ and $i\in\{1,\dots,M\}$. Therefore,
@@
By Proposition 1.3, there exists a similarity $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ such that $q\in f(K)\subset T$ and so that $f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$ is an open subset of $T$. Let $z\in f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$ and let $w\in A^{*}$ so that $f^{-1}(z)\in S_{w}(K)\subset(0,1)^{N}\cap K$. Then $f\circ S_{w}$ is a similarity so that $f\circ S_{w}((0,1)^{N}\cap K)$ is an open subset of $f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$, and therefore $f\circ S_{w}((0,1)^{N}\cap K)$ is an open subset of $T\cap B(q,\rho_{2})$. Therefore,
@@
is a similarity so that $(f\circ S_{w})\circ G^{-1}(B(p,\rho_{1})\cap T)$ is an open subset of $T\cap B(q,\rho_{2})$. ∎
@@
Let $s>1$, and given an even integer $n\geq 4$ recall numbers $\alpha_{n}$ from Section 5, sets $K^{n,0}$ from (8.1), and tangents $\{T_{n,k}\}_{k\in\mathbb{N}}$ of Proposition 8.1. Fix an even integer $n\geq 4$ so that $\alpha_{n}<s$ and a random choice function $\eta\in\mathscr{C}_{n}$. Let $I:K^{\eta}\to\Gamma_{s}$ be the snowflake map composed with the bi-Lipschitz embedding given in Section 9. Recall that $I$ is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism onto its image $\Gamma_{s}$, and $\Gamma_{s}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$. Furthermore, $\Gamma_{s}$ is a $(1/s)$-Hölder curve with $\mathcal{H}^{s}(\Gamma_{s})\in(0,\infty)$. Fix a point $x_{0}\in K^{\eta}$ so that for every $k\in\mathbb{N}$, $T_{n,k}\in\operatorname{Tan}(K^{\eta},x_{0})$, and recall by Lemma 9.1 for every $k\in\mathbb{N}$ there exists a tangent $T_{k}^{*}\in\operatorname{Tan}(\Gamma_{s},I(x_{0}))$ and an $(\alpha_{n}/s)$-bi-Hölder homeomorphism $I^{*}_{k}:T_{n,k}\to T^{*}_{k}$ induced by the map $I$. The next proposition completes the proof of Theorem 1.4.

### v3

@@ 8.1. Local cut-points in a class of carpets
To verify (i), fix such an open set $V\subset\mathbb{R}^{2}$. Then there exist $z\in V$ and $\epsilon>0$ such that $B(z,\epsilon)\subset V$ and $B(z,\epsilon/2)\cap L_{k}\neq\emptyset$. Assume for the sake of contradiction that (i) fails to hold for this $V$, and thus $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$ infinitely often. Now by Lemma 8.10 and since $\lim_{j\to\infty}Y_{N_{j}}=L_{k}$, there exists $j_{0}\in\mathbb{N}$ such that for every integer $j\geq j_{0}$, $\operatorname{excess}(Y_{N_{j}}\cap\overline{B}({\bf 0},|z|+\epsilon),X_{N_{j}})<\epsilon/2$ and $B(z,\epsilon/2)\cap Y_{N_{j}}\neq\emptyset$. By our assumption, there exists an integer $j\geq j_{0}$ such that $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$ and there exists a point $y\in Y_{N_{j}}\cap B(z,\epsilon/2)$. Then we have
@@
Now to verify (ii), fix $\delta>0$ and $z\in\mathbb{R}^{2}$ satisfying $B(z,\delta)\cap L_{k}=\emptyset$ and let $\epsilon\in(0,\delta)$. Since $Y_{N_{j}}\to L_{k}$ as $j\to\infty$ with respect to the Attouch-Wets topology, there is a $j_{0}\in\mathbb{N}$ such that for every integer $j\geq j_{0}$, $Y_{N_{j}}\cap B(z,\frac{\epsilon+\delta}{2})=\emptyset$. Furthermore by Lemma 8.10, choosing $j_{0}$ large enough, we may assume that for every integer $j\geq j_{0}$, $\operatorname{excess}(X_{N_{j}}\cap\overline{B}({\bf 0},|z|+\delta),Y_{N_{j}})<\frac{\delta-\epsilon}{2}$. Then for every integer $j\geq j_{0}$ and for each $p\in X_{N_{j}}\cap B({\bf 0},|z|+\delta)$, $\operatorname{dist}(p,Y_{N_{j}})<\frac{\delta-\epsilon}{2}$. Since $Y_{N_{j}}\cap B(z,\frac{\epsilon+\delta}{2})=\emptyset$, if $p\in B(z,\epsilon)$, then $\operatorname{dist}(p,Y_{N_{j}})\geq\frac{\delta-\epsilon}{2}$ and $p\in B({\bf 0},|z|+\delta)$. Therefore for each integer $j\geq j_{0}$, $X_{N_{j}}\cap B(z,\epsilon)=\emptyset$, as desired. ∎
@@
Assume first that $t\geq 1$. Let $x_{0}\in X$, let $T\in\operatorname{Tan}(X,x_{0})$, and let $r_{j}\to 0$ be a sequence of scales for which $r_{j}^{-1}(X-x_{0})\to T$ as $j\to\infty$ in the Attouch-Wets topology. For each $j\in\mathbb{N}$, define the function
@@
Let $R>0$, let $\epsilon>0$, and let $j_{0}\in\mathbb{N}$ such that for every integer $j\geq j_{0}$,
@@
The other excess estimate follows via a similar calculation, thus $r_{j}^{-1/t}(Y-f(x_{0}))\to F(T)$ as $j\to\infty$ in the Attouch-Wets topology. This concludes the case $t\geq 1$.
Now let $t\in(0,1)$. Then $f^{-1}:Y\to X$ is a $t$-bi-Hölder homeomorphism. Fix $x_{0}\in X$, let $T\in\operatorname{Tan}(X,x_{0})$, and let $r_{j}\to 0$ be a sequence of scales such that $r_{j}^{-1}(X-x_{0})\to T$ as $j\to\infty$ in the Attouch-Wets topology. Then $r_{j}^{t}\to 0$ is a sequence of scales, and there is a convergent subsequence $r_{j_{k}}^{-t}(Y-f(x_{0}))\to S$ as $k\to\infty$ in the Attouch-Wets topology. As in the case $t\geq 1$, we have $g_{j_{k}}:r_{j_{k}}^{-t}(Y-f(x_{0}))\to r_{j_{k}}^{-1}(X-x_{0})$, a sequence of $t$-bi-Hölder homeomorphisms with uniform coefficients which extend to $t$-Hölder functions $G_{j_{k}}:\mathbb{R}^{m}\to\mathbb{R}^{d}$, and the maps $G_{j_{k}}$ converge locally uniformly to a $t$-Hölder map $G:\mathbb{R}^{m}\to\mathbb{R}^{d}$. Then
@@
By Assouad’s embedding theorem [Ass83], there exists $N\in\mathbb{N}$ and a bi-Lipschitz embedding of the metric space $(K^{\eta},|\cdot|^{\alpha_{n}/s})$ into $\mathbb{R}^{N}$ and denote by $\Gamma$ the embedded image. This mapping produces an $(\alpha_{n}/s)$-bi-Hölder homeomorphism $g:K^{\eta}\to\Gamma$. Therefore, there exists some $C>1$ such that for every Borel $B\subset K^{\eta}$,
@@
so $\mathcal{H}^{s}(\Gamma)>0$. Moreover, the map $g\circ f:[0,1]\to\Gamma$ is a $(1/s)$-Hölder surjection. By Lemma 9.1, we have that for every point $x\in K^{\eta}$ and every tangent $T\in\operatorname{Tan}(K^{\eta},x)$ there exists a tangent $T^{\prime}\in\operatorname{Tan}(\Gamma,g(x))$ such that $T$ and $T^{\prime}$ are $(\alpha_{n}/s)$-bi-Hölder equivalent; in particular, $T$ and $T^{\prime}$ are homeomorphic to each other. Therefore, $\operatorname{Tan}(\Gamma,x)$ contains infinitely many topologically distinct elements. ∎
@@ 10. Tangents of self-similar sets
In this section we study tangents of self-similar sets. The main goals are the following proposition and the proof of Theorem 1.3.
Proposition 10.1.
Let $\{\phi_{i}:\mathbb{R}^{N}\to\mathbb{R}^{N}\}_{i=1}^{m}$ be an IFS of similarities satisfying the OSC for some open set $U$ with $U\cap Q\neq\emptyset$, where $Q$ is the attractor, and let $s=\operatorname{dim_{H}}{Q}$. There exists $C_{0}>1$ depending only on the Lipschitz norms of $\phi_{1},\dots,\phi_{m}$ such that for $\mathcal{H}^{s}$-almost every $x\in Q$, every $T\in\operatorname{Tan}(Q,x)$, and every $R>0$, there exist similarities $\{f_{i}:Q\to T\}_{i=1}^{n}$ with $n\leq C_{0}$ with Lipschitz norms in $C_{0}^{-1}R,R]$ such that
and for all distinct $i,j\in\{1,\dots,n\}$, $f_{i}^{R}(U)\cap f_{j}^{R}(U)=\emptyset$ and $f_{i}^{R}(U\cap Q)$ is an open subset of $T$. In particular, $\operatorname{dim_{H}}(T)=\operatorname{dim_{H}}(Q)$ and $\mathcal{H}^{\operatorname{dim_{H}}(Q)}(T)>0$.
Proof.
@@
Fix $x\in Q\cap U$, $T\in\operatorname{Tan}(Q,x)$, and $R>0$. Let $r_{j}\to 0$ be a decreasing sequence of positive scales such that $r_{1}<1/R$ and the sets $D_{j}:=(r_{j})^{-1}(Q-x)$ converge to $T$ in the Attouch-Wets topology as $j\to\infty$. Define for each $j\in\mathbb{N}$ the similarity $g_{j}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, given by $g_{j}(y)=(r_{j})^{-1}(y-x)$, and note that $g_{j}(Q)=D_{j}$. Define the alphabet $A=\{1,\dots,m\}$ and for a word $w\in A^{*}$, define $L_{w}:=\operatorname{Lip}(\phi_{w})$. Following [BV21, §2.3], for any $\delta\in(0,1)$ define
@@
Note that for each $j$, $\overline{B}(x,r_{j}R)\cap Q\subset\bigcup_{w\in C^{R}_{j}}\phi_{w}(Q)$. Additionally, by the OSC, if $v,w\in A^{*}$ are distinct words satisfying $\phi_{w}(U)\cap\phi_{v}(U)\neq\emptyset$, then there exists $u\in A^{*}\setminus\{\varepsilon\}$ such that either $w=vu$ or $v=wu$, which further implies that $\frac{L_{w}}{L_{v}}\leq L_{1}$ or $\frac{L_{v}}{L_{w}}\leq L_{1}$. Moreover, if $j\in\mathbb{N}$ and $w,v\in C^{R}_{j}$, then $\frac{L_{w}}{L_{v}}>L_{1}$ (or vice-versa); thus if $w,v\in C^{R}_{j}$ are distinct, $\phi_{w}(U)\cap\phi_{v}(U)=\emptyset$. Roughly speaking, each set $\{\phi_{w}:w\in C^{R}_{j}\}$ respects the “disjoint images” part of the open set condition for $U$. Observe that for each $j\in\mathbb{N}$ and each $w\in C^{R}_{j}$ by (10.1), the fact that $\operatorname{diam}(Q)=1$ and the fact that $\phi_{w}(Q)\cap\overline{B}(x,r_{j}R)\neq\emptyset$, we have $\phi_{w}(Q)\subset\overline{B}(x,2r_{j}R)\cap Q$. Hence, for each $j\in\mathbb{N}$,
@@
Next, since $Q$ is self-similar, it follows that $\mathcal{H}^{s}(Q)>0$ [Hut81, §5.3] and, in fact, $Q$ is Ahlfors $s$-regular [BV21, Lemma 2.4]. That is, there exists $c_{1}\geq 1$ such that for every $y\in Q$ and $r\in(0,1)$,
@@
Furthermore, if $\{f_{1}^{2R},\dots,f_{M_{2R}}^{2R}\}$ is constructed in a similar manner (replacing $R$ by $2R$), then for every $i\in\{1,\dots,M_{R}\}$ there exists unique $j\in\{1,\dots,M_{2R}\}$ such that $f_{i}^{R}(Q)\subset f_{j}^{2R}(Q)$. Indeed, the sets $C_{j}^{R}$ and $C_{j}^{2R}$ satisfy $\bigcup_{v\in C_{j}^{R}}\phi_{v}(Q)\subset\bigcup_{u\in C_{j}^{2R}}\phi_{u}(Q)$, thus for each $v\in C_{j}^{R}$, there exists a unique $u\in C_{j}^{2R}$ satisfying $\phi_{v}(U\cap Q)\subset\phi_{u}(U\cap Q)$. Then it must hold that $f_{i}^{R}(U\cap Q)$ is an open subset of $f_{j}^{2R}(U\cap Q)\cap B({\bf 0},2R)$ which is an open subset of $S\cap B({\bf 0},2R)$. Therefore, $f_{i}^{R}(U\cap Q)$ is an open subset of $S$. ∎
10.1. Tangents of self-similar sponges
@@
Suppose that $K$ is the attractor of $\{S_{i}\}_{i=1}^{m}$ with $S_{i}(y)=k^{-1}y+p_{i}$ where $k$ and $\{p_{1},\dots,p_{m}\}$ are as in (2.1). If $m=1$, then the result is trivial, hence we may assume that $m\geq 2$. Set $A=\{1,\dots,m\}$. Note that for each $n\in\mathbb{N}$, $v\in A^{n}$, and $i\in A$,
@@
If $(0,1)^{N}\cap K=\emptyset$, then the result is trivial, so assume $(0,1)^{N}\cap K\neq\emptyset$. Fix $x\in(0,1)^{N}\cap K$, $T\in\operatorname{Tan}(K,x)$, $p,q\in T$, $\rho_{1},\rho_{2}>0$, and a sequence of positive scales $r_{j}\to 0$ such that the sets $X_{j}:=(r_{j})^{-1}(K-x)$ converge to $T$ in the Attouch-Wets topology as $j\to\infty$.
@@
We claim that $\bigcap_{w\in W_{j}}S_{w}([0,1]^{N})\neq\emptyset$ for every $j\in\mathbb{N}$. We first show the claim for two words only. Fix distinct $v,w\in W_{j}$ and let $z_{v},z_{w}\in[0,1]^{N}$ such that
@@
thus by the triangle inequality we obtain
@@
By Lemma 2.9, for every $j\in\mathbb{N}$ there exists a face $C_{j}$ of $0,1]^{N}$ such that $\bigcap_{i=1}^{M}S_{v_{i}^{j}}([0,1]^{N})=S_{v_{1}^{j}}(C_{j})$. Thus, for every $j\in\mathbb{N}$ there exist distinct $e_{j,1},\dots,e_{j,M}\in\{0,1\}^{N}$ satisfying $S_{v_{i}^{j}}(e_{j,i})=S_{v_{1}^{j}}(e_{j,1})$ for every $i\in\{1,\dots,M\}$. Passing to a subsequence, we may assume that there exist distinct $e_{1},\dots,e_{M}\in\{0,1\}^{N}$ such that $e_{j,i}=e_{i}$ for every $j\in\mathbb{N}$ and $i\in\{1,\dots,M\}$. Then $f_{i}(e_{i})=f_{1}(e_{1})$ and $S_{v_{i}^{j}}(e_{i})=S_{v_{1}^{j}}(e_{1})$ for every $j\in\mathbb{N}$ and $i\in\{1,\dots,M\}$. Therefore,
@@
By Proposition 10.1, there exists a similarity $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ such that $q\in f(K)\subset T$ and such that $f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$ is an open subset of $T$. Let $z\in f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$ and let $w\in A^{*}$ satisfy $f^{-1}(z)\in S_{w}(K)\subset(0,1)^{N}\cap K$. Then $f\circ S_{w}$ is a similarity for which $f\circ S_{w}((0,1)^{N}\cap K)$ is an open subset of $f((0,1)^{N}\cap K)\cap B(q,\rho_{2})$, and therefore $f\circ S_{w}((0,1)^{N}\cap K)$ is an open subset of $T\cap B(q,\rho_{2})$. Thus,
@@
is a similarity such that $(f\circ S_{w})\circ G^{-1}(B(p,\rho_{1})\cap T)$ is an open subset of $T\cap B(q,\rho_{2})$. ∎
@@
Let $s>1$, and given an even integer $n\geq 4$ recall numbers $\alpha_{n}$ from Section 5, sets $K^{n,0}$ from (8.1), and tangents $\{T_{n,k}\}_{k\in\mathbb{N}}$ of Proposition 8.1. Fix an even integer $n\geq 4$ such that $\alpha_{n}<s$ and a random choice function $\eta\in\mathscr{C}_{n}$. Let $I:K^{\eta}\to\Gamma_{s}$ be the snowflake map composed with the bi-Lipschitz embedding given in Section 9. Recall that $I$ is an $(\alpha_{n}/s)$-bi-Hölder homeomorphism onto its image $\Gamma_{s}$, and $\Gamma_{s}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$. Furthermore, $\Gamma_{s}$ is a $(1/s)$-Hölder curve with $\mathcal{H}^{s}(\Gamma_{s})\in(0,\infty)$. Fix a point $x_{0}\in K^{\eta}$ such that for every $k\in\mathbb{N}$, $T_{n,k}\in\operatorname{Tan}(K^{\eta},x_{0})$, and recall by Lemma 9.1 for every $k\in\mathbb{N}$ there exists a tangent $T_{k}^{*}\in\operatorname{Tan}(\Gamma_{s},I(x_{0}))$ and an $(\alpha_{n}/s)$-bi-Hölder homeomorphism $I^{*}_{k}:T_{n,k}\to T^{*}_{k}$ induced by the map $I$. The next proposition completes the proof of Theorem 1.3.